### v1

@@ 4.2.2 Context Adaptability, Efficiency, and Transferability
Obs.❺ MCE is free from the inductive biases in context length imposed by prior methods. Prior CE methods exhibit inherent biases: GEPA [agrawal2025gepa] tends toward brevity, typically producing around 1–2K tokens, while ACE [zhang2025agentic] suffers from context bloat, reaching up to 80K tokens after 5 epochs of optimization (when using 200 training instances). In contrast, MCE learns to produce context of suitable length for each task. The two most effective contexts on FiNER have 1.5K and 20K tokens (Fig.˜3), while those on LawBench and USPTO50k reach 44K and 86K tokens. These results demonstrate that MCE adapts context length flexibly to task requirements, overcoming the inductive biases of prior methods.
@@ 5 Conclusion and Discussion
Future Work. We identify three meaningful directions for future research. First, the agentic skill evolution paradigm extends beyond CE. Prior evolutionary approaches target specific solutions, heuristic code, or search algorithms; MCE instead evolves skills, a higher-order and integrated abstraction essential for general AI. While static skills are well-recognized [anthropics_skills, muratcankoylan_agent_skills_context_engineering], MCE is among the first to dynamically evolve them, bridging manual skill engineering and autonomous self-improvement. The paradigm of evolutionary skills could generalize to other agentic capabilities and task domains. Second, our generator currently uses one-shot inference for consistency with CE baselines. Since MCE stores context as files, an agentic generator that interacts directly with these artifacts could be more effective. Extending MCE to co-evolve context utilization skills alongside context learning skills is a promising direction. Third, progressive disclosure is native to agent skills: agents load skill details into context only when relevant. This enables MCE to compose and scale skills across domains with minimal overhead. Investigating skill transfer across tasks and emergent behaviors from skill composition are interesting open questions.
@@ A.2 USPTO-50k
We utilize the USPTO-50k dataset [schneider2016s], a widely-used benchmark for single-step retrosynthesis prediction in computational chemistry. Given a target product molecule represented in SMILES notation, the task requires predicting the precursor reactants needed to synthesize it. This task evaluates the model’s understanding of organic reaction mechanisms and chemical transformation patterns.
We follow cai2025flex in data preprocessing. We randomly sample 50 instances from the original training split for training, 30 instances from the training split for validation, and 100 instances from the test split for evaluation. Stratified sampling is used to uniformly cover all 10 reaction types, ensuring balanced representation across different chemical transformation classes.
@@ A.5 Aegis2
We utilize the AEGIS2 dataset [ghosh2025aegis2], a comprehensive AI safety dataset designed for training lightweight guardrail models. Each instance contains a user prompt with its corresponding safety label (safe/unsafe) and violated risk categories. The task is binary classification of user prompts, with unsafe as the positive class and F1 score as the evaluation metric.
To identify challenging categories, we evaluate LLaMA-3-SafeGuard [dubey2024llama3herdmodels] across all 24 risk subcategories using the system prompts specified in the original paper. Based on per-category F1 scores, we selecte the four worst-performing categories: Copyright_Trademark_Plagiarism, Political_Misinformation_Conspiracy, Unauthorized_Advice, and Immoral_Unethical.

### v2

@@ 4.2.2 Context Adaptability, Efficiency, and Transferability
Obs.❺ MCE is free from the inductive biases in context length imposed by prior methods. Prior CE methods exhibit inherent biases: GEPA [Agrawal et al., 2025] tends toward brevity, typically producing around 1–2K tokens, while ACE [Zhang et al., 2026] suffers from context bloat, reaching up to 80K tokens after 5 epochs of optimization (when using 200 training instances). In contrast, MCE learns to produce context of suitable length for each task. The two most effective contexts on FiNER have 1.5K and 20K tokens (Fig.˜3), while those on LawBench and USPTO50k reach 44K and 86K tokens. These results demonstrate that MCE adapts context length flexibly to task requirements, overcoming the inductive biases of prior methods.
@@ 5 Conclusion and Discussion
Future Work. We identify three meaningful directions for future research. First, the agentic skill evolution paradigm extends beyond CE. Prior evolutionary approaches target specific solutions, heuristic code, or search algorithms; MCE instead evolves skills, a higher-order and integrated abstraction essential for general AI. While static skills are well-recognized [Anthropic, 2025b, Muratcan Koylan, 2025], MCE is among the first to dynamically evolve them, bridging manual skill engineering and autonomous self-improvement. The paradigm of evolutionary skills could generalize to other agentic capabilities and task domains. Second, our generator currently uses one-shot inference for consistency with CE baselines. Since MCE stores context as files, an agentic generator that interacts directly with these artifacts could be more effective. Extending MCE to co-evolve context utilization skills alongside context learning skills is a promising direction. Third, progressive disclosure is native to agent skills: agents load skill details into context only when relevant. This enables MCE to compose and scale skills across domains with minimal overhead. Investigating skill transfer across tasks and emergent behaviors from skill composition are interesting open questions.
@@ A.2 USPTO-50k
We utilize the USPTO-50k dataset [Schneider et al., 2016], a widely-used benchmark for single-step retrosynthesis prediction in computational chemistry. Given a target product molecule represented in SMILES notation, the task requires predicting the precursor reactants needed to synthesize it. This task evaluates the model’s understanding of organic reaction mechanisms and chemical transformation patterns.
We follow Cai et al. [2025b] in data preprocessing. We randomly sample 50 instances from the original training split for training, 30 instances from the training split for validation, and 100 instances from the test split for evaluation. Stratified sampling is used to uniformly cover all 10 reaction types, ensuring balanced representation across different chemical transformation classes.
@@ A.5 Aegis2
We utilize the AEGIS2 dataset [Ghosh et al., 2025], a comprehensive AI safety dataset designed for training lightweight guardrail models. Each instance contains a user prompt with its corresponding safety label (safe/unsafe) and violated risk categories. The task is binary classification of user prompts, with unsafe as the positive class and F1 score as the evaluation metric.
To identify challenging categories, we evaluate LLaMA-3-SafeGuard [Llama Team, 2024] across all 24 risk subcategories using the system prompts specified in the original paper. Based on per-category F1 scores, we selecte the four worst-performing categories: Copyright_Trademark_Plagiarism, Political_Misinformation_Conspiracy, Unauthorized_Advice, and Immoral_Unethical.